# Prompt-Enhanced Spatio-Temporal Graph Transfer Learning

### v1

@@ 4.2.1. Patching and Masking
As individual node signals contain low information density, reconstructing each signal separately could lead to the learning of trivial solutions (Shao et al., 2022). This issue can be more acute in the transfer setting when a target domain involves limited data. Therefore, we employ data patching (Liu et al., 2023e) to aggregate adjacent data into patches. Concretely, given sequential data $X_{i}$ from node $i$, we divide it into non-overlapping patches denoted by $\hat{X}_{i}\in\mathbb{R}^{T_{p}\times Ld_{x}}$, where $T_{p}$ is the number of patches and $L$ represents patch size. Then, we employ a linear projection to embed each patch into a feature space, represented as $S_{i}\in\mathbb{R}^{T_{p}\times d_{h}}$. Note that we will use $T$ to denote $T_{p}$ for simplicity in the following sections.
Next, we pre-train the model on source domains by randomly masking the patches. Contrasting with the graph or time series models that utilize noisy masking (Hou et al., 2022; Liu et al., 2023a), spatio-temporal random masking, which affects the two dimensions simultaneously, presents more complexity. As the Transformers only take in unmasked patches without padding (Shao et al., 2022), it is crucial to maintain a uniform patch number within each dimension to ensure a valid tensor structure, a requirement that noisy masking does not satisfy. Therefore, we divide the masking into spatial and temporal parts, where patches of a node or at a time step are all masked. Specifically, given a set of $N$ nodes and of $T$ steps, we randomly sample two sub-sets from them, denoted as $\mathcal{M}_{s}$ and $\mathcal{M}_{t}$. Then, a patch $s_{i}^{t}$ with its indexes in any of the sub-sets is marked as masked. The remaining unmasked patches, denoted by $S_{\bar{\mathcal{M}}}\in\mathbb{R}^{N_{u}\times T_{u}\times d_{h}}$ with ${S_{\bar{\mathcal{M}}}}=\{s_{i}^{t}|i\notin\mathcal{M}_{s}\text{ and }t\notin\mathcal{M}_{t}\}$, $N_{u}=N-|\mathcal{M}_{s}|$ and $T_{u}=T-|\mathcal{M}_{t}|$, are the inputs of the encoder. We set a total masking ratio of 75%, following (Shao et al., 2022; Hou et al., 2023). The scarcity of unmasked patches in the model aids in better capturing general spatio-temporal correlations.
@@ 4.2.2. Transformer-Based Encoder
Transformers have become dominant in pre-training frameworks due to their attractive learning capabilities (Khan et al., 2022; Kalyan et al., 2021; Wen et al., 2022). Additionally, masking further enhances their training efficiency on large-scale datasets by only feeding unmasked patches. Nevertheless, employing a single Transformer to model spatio-temporal dimensions still results in considerable time overhead, due to its complexity of $\mathcal{O}(N^{2}_{u}\times T^{2}_{u})$. To overcome this issue, we model the two dimensions respectively. We utilize TSFormer (Shao et al., 2022) to model temporal correlations and Graphormer (Ying et al., 2021) to capture spatial dependencies. Note that both models incorporate positional embeddings into their inputs. Following (Jin et al., 2023a), we enhance TSFormer by adding time of day (TOD) and day of week (DOW) embeddings, bolstering its capacity to capture periodic knowledge in the data-scarce target domain. In the bottom layer, representations learned from the two models are fused by a gating mechanism formulated as:
@@ 4.2.3. Gated Spatio-Temporal Decoder
The decoder first recovers full patches $H\in\mathbb{R}^{N\times T\times d_{h}}$ from $H_{\bar{\mathcal{M}}}$, where masked positions are replaced by a learnable mask token $h_{\mathcal{M}}$. The process is defined as:
@@
where $l$ represents a layer. In this way, the gating mechanism explicitly balances both spatial and temporal features in patches, facilitating the model on various tasks with different characteristics. Note that we still add the time embeddings at the first layer. Finally, we employ a linear prediction head to obtain patch predictions. We adopt mean absolute error (MAE) to optimize the model and calculate it solely on masked patches, in accordance with (He et al., 2022; Shao et al., 2022).
@@ 4.3. Two-Stage Prompting
In our transfer learning scenario, a pre-trained model needs to adapt to multiple downstream tasks on a target domain. Current prompting methods often introduce distinct solutions for domain transfer and task transfer (Ge et al., 2023; Sun et al., 2023). In contrast, we aim to achieve both goals within our single framework. Unfortunately, trivially optimizing a single bunch of prompts is inadequate for accomplishing the two goals. This is because domain transfer is designed to learn general domain knowledge, whereas task transfer focuses on capturing task-specific characteristics. Motivated by this observation, we introduce an innovative two-stage prompting pipeline that achieves these goals through a sequential, two-step approach.
@@ 4.4. STGP Learning Procedure
We describe the overall learning algorithm of STGP here. Given the encoder $f_{\theta_{e}}$, the decoder $f_{\theta_{d}}$, and the head $f_{\phi}$ parameterized by $\theta_{e}$, $\theta_{d}$, and $\phi$, we first train them through predictions of randomly masked patches. Then, keeping them frozen, the domain prompts $\pi_{dm}$ are optimized using the pre-training task for domain transfer. Finally, further with $\pi_{dm}$ held fixed, task prompts $\pi_{tk}$ are trained for a task. The head can optionally be tuned at this stage.
@@ 5.1. Experimental Settings
Baselines. We consider 18 baselines belonging to three categories.
@@
Evaluation Settings and Implementation We adopt the same setting as previous works (Lu et al., 2022; Liu et al., 2023e), where the datasets are divided into source, target, and test datasets. The source dataset comprises three datasets for pre-training. Then only the first three-day data in the remaining dataset are regarded as the target dataset for domain and task prompting. Finally, data from the remaining days serves as the test dataset, used for model evaluation. We adopt mean absolute error (MAE) and root mean squared error (RMSE) as the metrics. Hyperparameters are the same in all settings for STGP. Each patch, representing one-hour data, has a length of $L=12$, with $T=25$ patches in total. Each prompt bank contains 25 vectors. We run each experiment 5 times to report results.
@@ 5.3. Ablation Study
Figure 4. Model analysis with METR-LA as the target domain.
We conduct ablation studies by using METR-LA as the target domain. In each study, we modify the corresponding component while keeping other parts constant. For kriging and extrapolation, we adopt the inductive setting. The results are reported in Fig. 4.
@@ 6. Conclusion
In this paper, we present STGP, a novel prompt-based approach for spatio-temporal graph transfer learning. We develop a unified template alongside a task-agnostic network architecture suitable for multiple tasks. We then propose a two-stage prompting pipeline tailored for domain and task transfer. Comprehensive experiments and discussions demonstrate that STGP effectively captures domain knowledge and task-specific properties, consistently achieving state-of-the-art performance across diverse datasets and tasks.

### v2

@@ 4.2.1. Patching and Masking
As individual node signals contain low information density, reconstructing each signal separately could lead to the learning of trivial solutions (Shao et al., 2022). This issue can be more acute in the transfer setting when a target domain involves limited data. Therefore, we employ data patching (Liu et al., 2023e) to aggregate adjacent data into patches. Concretely, given sequential data $X_{i}$ from node $i$, we divide it into non-overlapping patches denoted by $\hat{X}_{i}\in\mathbb{R}^{T_{p}\times Ld_{x}}$, where $T_{p}$ is the number of patches and $L$ represents patch size. Then, we employ a linear projection to embed each patch into feature space, represented as $S_{i}\in\mathbb{R}^{T_{p}\times d_{h}}$. Note that we will use $T$ to denote $T_{p}$ for simplicity in the following sections.
Next, we pre-train the model on source domains by randomly masking the patches. Contrasting with the graph or time series models that utilize a noisy masking method (Hou et al., 2022; Liu et al., 2023a), spatio-temporal random masking, which affects the two dimensions simultaneously, presents more complexity. As the Transformers only take in unmasked patches without padding (Shao et al., 2022), it is crucial to maintain a uniform patch number within each dimension to ensure a valid tensor structure, a requirement that noisy masking does not satisfy. Therefore, we divide the masking into spatial and temporal parts, where patches of a node or at a time step are all masked. Specifically, given a set of $N$ nodes and of $T$ steps, we randomly sample two sub-sets from them, denoted as $\mathcal{M}_{s}$ and $\mathcal{M}_{t}$. Then, a patch $s_{i}^{t}$ with its indexes in any of the sub-sets is marked as masked. The remaining unmasked patches, denoted by $S_{\bar{\mathcal{M}}}\in\mathbb{R}^{N_{u}\times T_{u}\times d_{h}}$ with ${S_{\bar{\mathcal{M}}}}=\{s_{i}^{t}|i\notin\mathcal{M}_{s}\text{ and }t\notin\mathcal{M}_{t}\}$, $N_{u}=N-|\mathcal{M}_{s}|$ and $T_{u}=T-|\mathcal{M}_{t}|$, are the inputs of the encoder. We set a total masking ratio of 75%, following (Shao et al., 2022; Hou et al., 2023). The scarcity of unmasked patches in the model aids in better capturing general spatio-temporal correlations.
@@ 4.2.2. Transformer-Based Encoder
Transformers have become dominant in pre-training frameworks due to their attractive learning capabilities (Khan et al., 2022; Kalyan et al., 2021; Wen et al., 2022). Additionally, masking further enhances their training efficiency on large-scale datasets by only feeding in unmasked patches. Nevertheless, employing a single Transformer to model both spatial and temporal dimensions still results in considerable time overhead, due to its complexity of $\mathcal{O}(N^{2}_{u}\times T^{2}_{u})$. To overcome this issue, we model the two dimensions respectively by utilizing TSFormer (Shao et al., 2022) to model temporal correlations and Graphormer (Ying et al., 2021) to capture spatial dependencies. Note that both models incorporate positional embeddings into their inputs. Following (Jin et al., 2023a), we enhance TSFormer by adding time of day (TOD) and day of week (DOW) embeddings, bolstering its capacity to capture periodic knowledge in the data-scarce target domain. In the bottom layer, representations learned from the two models are fused by a gating mechanism formulated as:
@@ 4.2.3. Gated Spatio-Temporal Decoder
The decoder first recovers full patches $H\in\mathbb{R}^{N\times T\times d_{h}}$ from $H_{\bar{\mathcal{M}}}$, where masked positions are replaced by a shared learnable mask token $h_{\mathcal{M}}$:
@@
where $l$ represents the layer index. In this way, the gating mechanism dynamically balances spatial and temporal patterns, facilitating the model on various tasks with different characteristics. Note that we still add the time embeddings at its first layer. Finally, we employ a linear prediction head to obtain patch predictions. We adopt mean absolute error (MAE) to optimize the model and calculate it solely on masked patches, in accordance with (He et al., 2022; Shao et al., 2022).
@@ 4.3. Two-Stage Prompting
In our transfer learning scenario, a pre-trained model needs to adapt to multiple downstream tasks in a target domain. Current prompting methods often introduce distinct solutions for domain transfer and task transfer (Ge et al., 2023; Sun et al., 2023). In contrast, we aim to achieve both goals within our single framework. Unfortunately, trivially optimizing a single bunch of prompts is inadequate for accomplishing the two goals. This is because domain transfer is designed to learn general domain knowledge, whereas task transfer focuses on capturing task-specific characteristics. Motivated by this observation, we introduce an innovative two-stage prompting pipeline that achieves these goals through a sequential, two-step approach.
@@ 4.4. STGP Learning Procedure
We present the overall learning algorithm of STGP in Algo. 1. Given the encoder $f_{\theta_{e}}$, the decoder $f_{\theta_{d}}$, and the head $f_{\phi}$ parameterized by $\theta_{e}$, $\theta_{d}$, and $\phi$, we first train them through predictions of randomly masked patches. Then, keeping them frozen, the domain prompts $\pi_{dm}$ are optimized using the pre-training task for domain transfer. Finally, further with $\pi_{dm}$ held fixed, task prompts $\pi_{tk}$ are trained for a task. The head can optionally be tuned at this stage.
Algorithm 1 STGP Pre-Training and Prompting
0:  Source and target domain data $\mathcal{G}^{\mathcal{S}}$, $\mathcal{G}^{\mathcal{T}}$. Encoder $f_{\theta_{e}}$, decoder $f_{\theta_{d}}$, head $f_{\phi}$. Domain and task prompts $\pi_{dm}$ and $\pi_{tk}$.
0:  Optimal parameters $\theta_{e}$, $\theta_{d}$, $\phi$, $\pi_{dm}$ and $\pi_{tk}$.
1:  // Pre-training $f_{\theta_{e}}$, $f_{\theta_{d}}$ and $f_{\phi}$ with $\mathcal{G}^{\mathcal{S}}$
2:  while not converged do
3:     $S_{\bar{\mathcal{M}}},S_{\mathcal{M}}\longleftarrow\text{ PatchingMasking }(\mathcal{G}^{\mathcal{S}})$
4:     $\hat{S}_{\mathcal{M}}=f_{\phi,\theta_{e},\theta_{d}}(S_{\bar{\mathcal{M}}})$triangleright  Data reconstruction, Eq. 46
5:     Calculate loss $\operatorname{MAE}(\hat{S}_{\mathcal{M}},S_{\mathcal{M}})$ and optimize $\theta_{e}$, $\theta_{d}$, and $\phi$.
6:  // Domain prompting using $\mathcal{G}^{\mathcal{T}}$
7:  $f_{\theta_{e}},f_{\theta_{d}},f_{\phi}\longleftarrow$ Freeze $(\theta_{e},\theta_{d},\phi)$
8:  while not converged do
9:     $S_{\bar{\mathcal{M}}},S_{\mathcal{M}}\longleftarrow\text{ PatchingMasking }(\mathcal{G}^{\mathcal{T}})$
10:     $S_{\bar{\mathcal{M}}}^{*}\longleftarrow$ DomainPrompting $(S_{\bar{\mathcal{M}}},\pi_{dm})$triangleright  Eq. 8
11:     $\hat{S}_{\mathcal{M}}=f_{\phi,\theta_{e},\theta_{d}}(S_{\bar{\mathcal{M}}}^{*})$triangleright  Eq. 46
12:     Calculate loss and optimize $\pi_{dm}$.
13:  // Task prompting
14:  $f_{\theta_{e}},f_{\theta_{d}},f_{\phi},\pi_{dm}\longleftarrow$ Freeze $(\theta_{e},\theta_{d},\phi,\pi_{dm})$
15:  while not converged do
16:     $S_{\bar{\mathcal{M}}},S_{\mathcal{M}}\longleftarrow\text{ PatchingMasking }(\mathcal{G}^{\mathcal{T}})$triangleright  Downstream task masking
17:     $S_{\bar{\mathcal{M}}}^{*}\longleftarrow$ DomainPrompting $(S_{\bar{\mathcal{M}}},\pi_{dm})$triangleright  Eq. 8
18:     $H_{\bar{\mathcal{M}}}=f_{\theta_{e}}(S^{*}_{\bar{\mathcal{M}}})$triangleright  Eq. 4
19:     $H\longleftarrow$ ReocverFullPatches $(H_{\bar{\mathcal{M}}})$triangleright  Eq. 5
20:     $H^{*}\longleftarrow$ TaskPrompting $(H,\pi_{tk})$triangleright  Eq. 9
21:     $\hat{S}_{\mathcal{M}}=f_{\phi,\theta_{d}}(H^{*})$triangleright  Eq. 6
22:     Calculate loss and optimize $\pi_{tk}$triangleright  $\phi$ can be tuned optionally
23:  return  Trained models $f_{\theta_{e}},f_{\theta_{d}},f_{\phi}$ and prompts $\pi_{dm},\pi_{tk}$
@@ 5.1. Experimental Settings
Baselines. We consider 19 baselines belonging to three categories.
@@
Evaluation Settings and Implementation. We adopt the same setting as previous works (Lu et al., 2022; Liu et al., 2023e), where the datasets are divided into source, target, and test datasets. The source datasets comprise three datasets for pre-training. We sequentially split them into training and validation segments with a ratio of $8:2$, where the model is selected on the validation part. Then only the first three-day data in the remaining dataset is regarded as the target dataset for domain and task prompting. The subsequent three-day data is utilized for model selection and the rest of the segment is adopted to compute metrics. We adopt mean absolute error (MAE) and root mean squared error (RMSE) as the metrics.
Hyperparameters are the same in all settings for STGP. Each patch, representing one-hour data, has a length of $L=12$, with $T=25$ patches in total. Both Graphormer and TSFormer contain 4 layers with a hidden size of 128. The decoder contains 6 spatio-temporal gated layers with a hidden size of 32. We also add skip and residual connections into the decoder. The prediction head is a two-layer linear projection with a size of 128 and 256. All prompt banks involve 25 vectors with a size of 128. For extrapolation, we fine-tune the prediction head when optimizing the task prompts, as we empirically find it can yield better results. We run each experiment 5 times to report results. Our model is implemented by PyTorch and trained on one NVIDIA A100 GPU with 32GB memory. The source code is available at https://github.com/hjf1997/STGP.
@@ 5.3. Ablation Study
We conduct ablation studies on the three tasks by using METR-LA as the target domain. For kriging and extrapolation, we adopt the inductive setting. The results are reported in Fig. 4.
@@
Figure 4. Model analysis with METR-LA as the target domain.
@@ 6. Conclusion
In this paper, we present STGP, a novel prompt-based approach for spatio-temporal graph transfer learning. We develop a unified template alongside a task-agnostic network architecture suitable for multiple tasks. We then propose a two-stage prompting pipeline tailored for domain and task transfer. Comprehensive experiments and discussions demonstrate that STGP effectively captures domain knowledge and task-specific properties, consistently achieving state-of-the-art performance across diverse datasets and tasks. Regarding future work, we find that the downstream tasks unified by STGP fall under predictive learning, where known data is utilized to predict unknown data in a regression setting. Although received less attention, spatio-temporal graphs also encompass other applications, such as classification. Exploring the applicability of our unified template and network architecture to these tasks presents an intriguing direction for future research.
Acknowledgements.
This research is supported by Singapore Ministry of Education Academic Research Fund Tier 2 under MOE’s official grant number T2EP20221-0023. It is also part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.